UARS Weekly Status Report

01 February 2002




********************************************************************

                UARS WEEKLY STATUS REPORT - 01 FEB 2002

********************************************************************

***********
SOLSTICE
***********

Nothing received.

***********
SUSIM
***********

Nothing received.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

     PEM has also noticed the effect reported by HALOE that the UARS Orbit and 
Attitude Services (UOAS) software accesses the current processing day 
file but bombs when accessing the next day file.  Processing two consecutive 
files is normal for PEM, which process orbit and attitude data across the day 
(file) boundary.  This error occurs on CDHFB:: and was not previously noticed
on CDHF2:: or when Production was processing PEM data.  PEM has not checked
UOAS operation on CDHF2:: as yet.

     An initial review of the PEM CICEL PROCs were conducted.  Comments have 
been submitted.  The converted STOL procedures are not acceptable for use at 
the current time.  PEM is currently waiting for corrected versions of the 
PEM CICEL PROC to conduct further reviews.

     PEM continues to receive NMC data files from NOAA.  No problems have
been encountered.

     PEM awaits support of the UARS Traceability Mission for the TIMED mission
and has requested meetings with a GUVI team representative.


*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

The MPG participated in the UARS Ground System
Rehost Meeting on 28 January 2002.

The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop 
capabilities for the Science Traceability Mission.

Daily processing of instrument Daily Activity Plans 
continues with no problems.

***********
CDHF
***********

Nothing received.
                  
***********
FOT
***********

                           REPORTING PERIOD

This  report covers 19 January 2002 (Orbit 56628, GMT Day 019) through
25 January 2002 (Orbit 567322, GMT Day 025).

                         SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and   is  performing  nominally  under  two-battery  operations.   The
instrument operational changes were:


                    INSTRUMENT OPERATIONAL CHANGES


       Instrument    Time          Orbit     Comment
       PEM           021/2330      56672     AXIS-2 POWER OFF
       HALOE         022/1238      56680     Power ON


The  beta  angle ranged between the angles of 68.19 degrees  to  39.08
degrees  for  this  report period.  The beta angle is  now  decreasing
towards a minimum angle of 0 degrees on 03 February 2002 (DOY 034).
   
Spacecraft  battery  1,  2,  and 3 performance  monitoring  continues.
Battery 1 remains OFF the charge relay.  Battery 2 and Battery  3  are
currently maintaining greater than 25.44 V end-of-night (EON) load bus
voltage.   Battery  temperatures are stable with a  temperature  delta
between Battery 2 and Battery 3 of 1.25 to 1.31 degrees C.  Additional
battery performance data for each battery on 22 December 2001  and  28
December 2001 are provided in Appendix D.

The  clock  error  ranged from 9.5 to 11.3 msec  during  this  report
period.  Clock rate adjustments are listed in Appendix C.

All  instrument  and subsystem engineers were advised on  the  initial
excursion  of ALERTS or Out-Of-Limits conditions detected by  the  FOT
for  this report.  A summary is listed below and the details of  ALERT
and Out-Of-Limits occurrences are listed in Appendix B.


    OUT-OF-LIMITS  =    00

    ALERTS         =    00

The  following real-time command sheets were executed for  the  listed
operational element(s) during this report period (see Appendix A).


    FOT            =    2

    PEM            =    2

    HALOE          =    3

UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.


    FOT            =    0

One (1) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).

TRANSPONDER FREQUENCIES:
    Transponder A:  01/333  - 2287.496855
    Transponder B:  01/324  - 2287.498688


                        GROUND SYSTEM OPERATION


Zero  (0) operational support problems occurred this period.   Anomaly
Reports are included in Appendix E.

The  data  loss  calculations are no longer  being  supplied  by  Data
Capture Facility (DCF). This service should resume soon.



                       OTHER SIGNIFICANT EVENTS


                       UARS SATELLITE OPERATIONS


The  FOT received the new delivery of the Eclipse software as expected
on  25 January 2002.  This new release includes the loading capability
and   should   hopefully  resolve  the  remaining  telemetry   issues.
Verifying  these  capabilities will be the focal  point  of  simulator
testing this week.  The load testing needs to be completed in order to
keep with the scheduled first load to the spacecraft next week.


                               MEETINGS

The  FOT  attended  a  UARS Ground System Rehost  Meeting  on  Monday,
January 21, 2002.

The  FOT  attended a UARS Eclipse Working Meeting on Thursday, January
24, 2002.


                       FUTURE SIGNIFICANT EVENTS


The  next  UARS Yaw slew is scheduled for 03 February 2002 (DOY  034).
This  maneuver will be a Reverse to Forward Flight maneuver  (Southern
Hemisphere viewing).


Attachments:    Appendix A, Real-time Command Sheets
                Appendix B, Out-Of-Limits & ALERT Occurrences
                Appendix C, Clock & Solar Array Rate Corrections
                Appendix D, UARS Battery Performance Data
                Appendix E, UARS Anomaly Reports
                              APPENDIX A
                       REAL-TIME COMMAND SHEETS
                      Week ending 25 January 2002

ORBIT   AOS        NO.      SUBSYS      REASON
-----   ---        ---      ------      ------
56672   021/2320   02-019   PEM         ZEPS/HEPS  to DAY-ONLY OPS
56672   021/2330   02-020   PEM         AXIS-2 POWER OFF
56679   022/0939   02-021   OBC         Clk Adj  20.6 to 0.0 msec/day
56680   022/1238   02-022   HALOE       Power ON
56680   022/1238   02-023   HALOE       AZIMUTH FIX
56696   023/1313   02-024   HALOE       STOW POSITION FIX
56715   024/1918   02-025   OBC         Clk Adj 0.0 to 20.6 msec/day

                                   
                              APPENDIX B
                  OUT-OF-LIMITS & ALERTS OCCURRENCES
                      Week ending 25 January 2002

The following Out-Of-Limits occurrences were detected:

MNEMONIC       STATE   DESCRIPTION                ORBITS
--------       -----   -----------                ------
-None this report period.


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION                          ORBITS
--------     -----------                          ------
-None this report period.

_____________________________________________________________________
                              APPENDIX C
                 CLOCK & SOLAR ARRAY RATE CORRECTIONS
                      Week ending 25 January 2002

                     CLOCK DRIFT RATE ADJUSTMENTS

ORBIT          TIME                   CLOCK RATE CHANGE
-----          -----                  -----------------
56679          022/0939               Clk Adj  20.6 to 0.0 msec/day
56715          024/1918               Clk Adj 0.0 to 20.6 msec/day


                              APPENDIX D
                     UARS BATTERY PERFORMANCE DATA
                      Week ending 25 January 2002


  19 January 2002 - GMT Day 019 Beta = 68.19 deg, SA Pos = 269 deg

     DIFF V (mV)  TEMP (C)    CURR (amp)   EON V   AVE C/D  AVE
BAT  MAX/MIN      MAX/MIN     MAX/MIN      MIN     RATIO    DOD(%)
---  -------      --------    ----------   -----   -------  ------
1  -560.0/-728.0 -2.85/-3.77  +0.0/ -0.0   10.72    0.000    0.00
2   000.0/-224.0 +3.82/+2.52  +18.4/-13.2  27.04    1.040   15.15
3  +224.0/-224.0 +2.85/+1.88  +17.6/-12.8  27.04    1.040   14.91


  25 January 2002 - GMT Day 025 Beta = 39.08 deg, SA Pos = 269 deg


     DIFF V (mV)  TEMP (C)    CURR (amp)   EON V   AVE C/D  AVE
BAT  MAX/MIN      MAX/MIN     MAX/MIN      MIN     RATIO    DOD(%)
---  -------      --------    ---------    -----   -------  ------
1  -504.0/-672.0 -3.77/-4.69  +0.0/ -0.0   10.72    0.000    0.00
2  +112.0/-448.0 +5.14/+4.48  +28.0/-14.4  25.76    1.030   20.74
3  +224.0/-112.0 +4.15/+3.17  +27.2/-14.4  25.44    1.020   20.39

      V/T LEVEL CHANGES
ORB#          TIME       LVL FR-TO
-----         ----       ---------
-no changes    V/T 5

   





                              APPENDIX E
                         UARS ANOMALY REPORTS
                      Week ending 25 January 2002

                     NEW ANOMALY REPORTS GENERATED
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
-None this report period.

                ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
-None this report period.

                  ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
-None this report period.

                         OPEN ANOMALY REPORTS
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
02-002    FOT       see list  Jan 02    14170     Generic Late Acquisition